OpenStudy (anonymous):

the formula for the cylindrical volume of an engine on a dragster is (pi/4)B^2s, where B is the inside diameter (the bore) and S is the distance that the piston moves from its highest position to its lowest position (the stroke). Following the order of operations, describe the steps you would take to simplify the formula.

OpenStudy (agent0smith):

\[\Large \left( \frac{ \pi }{ 4 } \right)B^2 S\]like this?

OpenStudy (anonymous):

yea

OpenStudy (agent0smith):

I guess they want something like Divide pi by 4, then multiply that by the square of B, and then multiply by S

OpenStudy (anonymous):

what that is so confusing Dan

OpenStudy (agent0smith):

Idk how else to explain it...

OpenStudy (anonymous):

im not that smart lol :P

OpenStudy (agent0smith):

lol there's no other way to explain the steps Divide pi by 4, then multiply by by the square of B, then multiply by S

OpenStudy (anonymous):

12.57 =pi/4

OpenStudy (agent0smith):

that looks like pi times 4. Not pi divided by 4

OpenStudy (anonymous):

ooops lol

OpenStudy (anonymous):

my bad *blushes* :)

OpenStudy (agent0smith):

haha it's okay

OpenStudy (anonymous):

i feel so stupid now lmfao

OpenStudy (agent0smith):

lol it's okay!

OpenStudy (anonymous):

ok *still blushes in corner*

OpenStudy (agent0smith):

idk how else to help... it says "Following the order of operations, describe the steps you would take to simplify the formula." i think your answer is just this Divide pi by 4, then multiply that by the square of B, and then multiply by S thats all you need

OpenStudy (anonymous):

ok thanks Dan